# Collective excitations in chiral spin liquid: chiral roton and long-wavelength nematic mode

### v1

@@ Methods
where $|0\rangle$ is the ground-state wavefunction. To obtain the DSF of each local operator $\hat{O}$, In practice, we evolve the state $\psi(t)=e^{-\mathrm{i}\hat{H}t}\hat{O}_{\mathbf{0}}|0\rangle$ step by step with $\Delta t=0.05$, and the correlator is measured after each time step. For the results in the main text, we evolve totally $N_{t}=1000$ steps such that the TDVP spectra are obtained on a frequency grid with spacing $\Delta\omega=0.02$. In the time-to-frequency Fourier transform, we have used the Lorentzian broadening $\eta=0.05$. The bond dimension for the TDVP results is up to $\chi=400$, which provides well converged results for the studied systems in this work sup . TeNPy library is used in these matrix-product-state simulations Hauschild et al. (2024).
@@
Author contributions
H.L. initiated the project and performed the numerical simulations. All authors contributed to the discussion, analysis of the data and writing of the manuscript.
Data availability
The data supporting the plots within this paper and other findings of this study are available from the corresponding author upon reasonable request.
@@ Acknowledgements.
Acknowledgments
H. L. thanks Jie Wang and Xin Shen for helpful discussions, and Xiao-Tian Zhang and Han-Qing Wu for some ground-state benchmark, and Min Long, Zeno Bacciconi, Hernan B. Xavier, Zi Yang Meng, and Marcello Dalmonte for collaborations in studying the FCI excitations, and Zheng Yan for the hospitality at the early stage of this project. W. Z. thanks S. S. Gong, J. W. Yang, L. Wang for collaborating and fruitful discussion on the previous projects. H. L. and W. Y. are supported by the National Natural Science Foundation of China (No. 12425406), Research Grant Council of Hong Kong (AoE/P-701/20, HKU SRFS21227S05), and New Cornerstone Science Foundation. We thank Beijing PARATERA Tech Co., Ltd. (https://cloud.paratera.com) for providing HPC resources that supported the research results reported in this paper. This work is also supported by the National Key Research and Development Program of China Grant No. 2022YFA1402204 .
@@ Acknowledgments
In Fig.4 of the main text, we have shown the DSFs in the chiral p-wave channels with fixed $\omega=0.28$ (the energy scale of the roton minimum). Here, as a complement, we show their $\omega$-dependence in Fig.S5. The dispersion of the chiral roton is the same as that from the $\hat{B}_{\mu}$ operators and the spectral response fades away at other momenta.

### v2

@@ Methods
where $|0\rangle$ is the ground-state wavefunction. To obtain the DSF of each local operator $\hat{O}$, In practice, we evolve the state $\psi(t)=e^{-\mathrm{i}\hat{H}t}\hat{O}_{\mathbf{0}}|0\rangle$ step by step with $\Delta t=0.05$, and the correlator is measured after each time step. For the results in the main text, we evolve totally $N_{t}=1000$ steps such that the TDVP spectra are obtained on a frequency grid with spacing $\Delta\omega=0.04\pi$. In the time-to-frequency Fourier transform, we have used the Lorentzian broadening $\eta=0.05$. The bond dimension for the TDVP results is up to $\chi=400$, which provides well converged results for the studied systems in this work sup . TeNPy library is used in these matrix-product-state simulations Hauschild et al. (2024).
@@
Supplemental Information
The supplemental information contains:
I. the ground-state entanglement spectrum,
II. additional ED and TDVP results of the collective modes,
III. additional information of TDVP simulations,
to support the discussion in the main text.
@@ Acknowledgements.
Acknowledgments —  H. L. thanks Jie Wang and Xin Shen for helpful discussions, and Xiao-Tian Zhang and Han-Qing Wu for some ground-state benchmark, and Min Long, Zeno Bacciconi, Hernan B. Xavier, Zi Yang Meng, and Marcello Dalmonte for collaborations in studying the FCI excitations, and Zheng Yan for the hospitality at the early stage of this project. W. Z. thanks S. S. Gong, J. W. Yang, L. Wang for collaborating and fruitful discussion on the previous projects. H. L. and W. Y. are supported by the National Natural Science Foundation of China (No. 12425406), Research Grant Council of Hong Kong (AoE/P-701/20, HKU SRFS21227S05), and New Cornerstone Science Foundation. We thank Beijing PARATERA Tech Co., Ltd. (https://cloud.paratera.com) for providing HPC resources that supported the research results reported in this paper. This work is also supported by the National Key Research and Development Program of China Grant No. 2022YFA1402204 .
@@
In Fig.4 of the main text, we have shown the DSFs in the chiral p-wave channels with fixed $\omega=1.76$ (the energy scale of the roton minimum). Here, as a complement, we show their $\omega$-dependence in Fig.S5. The dispersion of the chiral roton is the same as that from the $\hat{B}_{\mu}$ operators and the spectral response fades away at other momenta.